SpaceRef

SpaceRef


NASA Space Station On-Orbit Status 26 May 2005

Status Report From: NASA HQ/SpaceRef/NASA Watch
Posted: Thursday, May 26, 2005

image SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2005) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.

FE/SO John Phillips set up the ADUM (Advanced Diagnostic Ultrasound in Micro-G) equipment, including video and electronic still cameras, and later performed the first cardiac /thoracic (Scan A) investigation for Increment 11, with CDR Krikalev as subject. Due to today's Ku-band comm pattern, the session was split into three different parts. The MGS (Microgravity Science Glovebox) video camcorders were used in lieu of the Lab VTRs, requiring prior activation of the MSG and subsequent video tape replacement, before closeout ops. [With Sergei strapped down on the CMRS (crew medical restraint system) and wearing electrodes for ECG (electrocardiogram) recording, John performed the ultrasound scans, unassisted by the subject in order to investigate scans on an incapacitated subject. Afterwards the hardware was deactivated and the scan heads were cleaned and stowed.]

After deactivating two fans (VPO5, VPO6) behind Service Module (SM) panel 307 (Zone III Overhead), Krikalev installed modified mufflers on the two fan housings, afterwards activating the fans and closing the panel. [This installation had to be aborted on 5/23, when the insulation exceeded available space by about 2 cm, hindering panel closure without compressing the muffler textile material.]

Phillips performed the bi-monthly maintenance reboot of the OCA comm SSC (Station Support Computer) router laptop.

Shuttle
Processing Status
News
Daily Mission
Return to Flight
ISS
Weekly Status
Weekly Science
Daily On-Orbit Status
Daily Crew Timeline
Soyuz | Progress
ISS News | ATV

John also conducted the daily routine maintenance of the SM's SOZh environment control & life support system, including its ASU toilet system plus today the BRPK air/liquid condensate separator apparatus, and prepared the regular IMS "delta" file for export/import to the IMS databases.

Working off his voluntary "job jar" task list, the CDR conducted another session with the "Uragan" (hurricane) earth-imaging program, focusing the Nikon D1X No. 3 digital camera with 800-mm lens on targets called out on an uplinked list. [Targets for today included such views as the Katun River Valley, Lake Teletskoye, the Sayan Mountains and the city of Kyzyl, forests on the S bank of the Belaya River at the confluence with the Angara, panoramic imagery of Lake Baikal, the Ussuri River flood, the city of Adler, the Mzymta River and the skiing resort near Krasnaya Poliana, Krasnodar Territory, areas west and north of the city of Astrakhan (real-time downlink), the southern section of the Volga Delta adjacent to the Caspian Sea, and the city of Guriev near the Ural River.]

Both crewmembers conducted their regular 2.5-hr. physical exercise program on the TVIS treadmill (CDR), CEVIS cycle ergometer (FE), RED resistive machine and VELO bike with bungee cord load trainer. [TVIS could not be scheduled for John today due to conflicts with other activities in the SM but will again be timelined on Friday. Sergei's daily protocol prescribes a strict four-day microcycle exercise with 1.5 hr on the treadmill and one hour on VELO plus load trainer (today: Day 3 of a new set).]

Afterwards, the FE transferred the exercise data files to the MEC for downlink, as well as the daily wristband HRM (heart rate monitor) data of workouts on CEVIS and RED, followed by their erasure on the HRM storage medium.

Yesterday's CONUS VHF (very-high frequency) proficiency comm pass over NASA VHF sites at Dryden and White Sands (WSC) went well, except for some severe noise interference during the WSC pass. The problem is under investigation at WSC.

The automated charge/discharge maintenance cycle on the two EMU batteries #2047 & #2048 in the US Airlock (A/L) continues to go well. [Charging took about 16 hours, followed by one hour cooling and then another 16 hours of discharging at present.]

The partially depleted PBA (Portable Breathing Apparatus) oxygen bottle found during the 5/24 inspection at the Node Port 2 location was determined to have been accidentally discharged when restowed after an earlier inspection. Yesterday, the crew completely discharged the O2 bottle into the cabin, stowed it, and replaced it with a fresh unit from the A/L (leaving two filled bottles in the A/L).

To reestablish condensate flow for the currently stalled SKV air conditioner system, TsUP/Moscow is planning for the crew to replace the sediment-clogged F3 connector at the SRVK-2M (condensate water recovery system) inlet valve, i.e., SKV/SRVK interface, as was done before. Meanwhile, the CCAA (Common Cabin Air Assembly) air conditioner in the Lab is collecting condensate. [Plans are to remove another connector (K27) downstream of F3, clean the residue out, and replace it.]

The first step of the software transition of the Russian computer system to Vers. 7.03 yesterday went well, as reported, except for a missing cable connection between Central Post computer 2 ( KTsP-2) and Laptop 2 (LT2), which temporarily prevented the latter's upgrading. [The next steps of the transition will be performed on 5/29-6/1, consisting of first activation of the SM Central (TsVM) & Terminal (TVM) Computers with the new load, checkout of the TsVM/TVM connection with a control subroutine plus telemetry monitoring during RGS (Russian ground site), and completion of the upgrade, followed by operational activation. By 6/1, the transition should be complete.]

Yesterday's attempt by Krikalev to install a network channel controller box (KSK A24) for ATV PCE (Automated Transfer Vehicle Proximity Communication Equipment) testing and hook it up to the BITS2-12 onboard telemetry measurement system had to be aborted when two connectors could not be mated. TsUP is developing a resolution plan. Reinstallation of the temporarily removed VPO12 fan in the SM has also been deferred.

More diagnostic work on the currently defunct Elektron O2 generator is planned for tomorrow. [The electrical continuity testing with the MMTs-01 Volt/Amp Multimeter and a special pin socket adapter will not be performed on the Liquid Unit (BZh) itself, but on the suspect Signal & Command Matching Unit (BSSK) between the Elektron electrolysis machine and the station's BITS2-12 system (the BSSK was replaced last on 1/22/04). A new BSSK will be manifested next month on Progress 18, for potential replacement.]

Update on Solid-Fuel Oxygen Generator (SFOG) "candles": As of tonight, since 5/20 a total of 9 candles have been decomposed on board (total attempts: 10 [i.e., 1 dud]).

Today's CEO (crew earth observations) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, were Caracas, Venezuela (this pass, in good light, was early enough in the day to prior to significant cloud formation. The Venezuelan capital city lies in a valley just inland from the coast. Shooting just left of track, with the long lens for detail here), Central-Arizona Phoenix (while this pass was over the extreme NW corner of this target area, the crew was to take advantage of the sun glint enhancement of the irrigated agriculture in the Gila Valley between Yuma and Phoenix), and Andrews Forest, Washington (ISS had a nice pass with good illumination over the southern portion of the area of this Long Term Environmental Research site [LTER, see 5/14 Status report]).

CEO photography can be viewed and studied at the websites:

See also the website "Space Station Challenge" at:

To view the latest photos taken by the expedition 11 crew visit:

Expedition 11 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

ISS Location NOW

Full Size/Update
Real Time ISS Tracker - More Links

ISS Orbit (as of this morning, 6:37am EDT [= epoch]):

  • Mean altitude -- 352.8 km
  • Apogee height -- 356.4 km
  • Perigee height -- 349.2 km
  • Period -- 91.60 min.
  • Inclination (to Equator) -- 51.65 deg
  • Eccentricity -- 0.0005398
  • Solar Beta Angle -- -19.4 deg (magnitude increasing)
  • Orbits per 24-hr. day -- 15.72
  • Mean altitude loss in last 24 hours -- 52 m
  • Revolutions since FGB/Zarya launch (Nov. 98) -- 37226

Some Increment 11 Main Events (not final):

  • __Progress M-52 (17P) undock -- 6/16;
  • __Progress M-53 (18P) launch -- 6/17 (dock 6/19);
  • __Reboost -- ~6/22 (delta-V 1.5 m/s);
  • __LF-1/STS-114 launch -- 7/13 (8-day window opens);
  • __Soyuz TMA-6 (10S) relocate (from DC-1 to FGB) -- ~8/16;
  • __ULF1.1/STS-121 launch -- NET 9/9 (8-day window opens);
  • __12A/STS-115 launch -- NET 2/16/06;
  • __12A.1/STS-116 launch -- NET 4/23/06;
  • __13A/STS-117 launch -- NET 5/18/06.

ISS Altitude History

Apogee height -- Mean Altitude -- Perigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA's Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA's Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.